OpenStudy (anonymous):

@amistre64 Can you help me Literal Equation Solve S = 4v2 for v.

OpenStudy (lacypennelll):

Divide by 4 to get s/4=v^2 Take the square root to get v=+/- sqrt(s)/2

OpenStudy (anonymous):

Can i get a visual of that?

OpenStudy (anonymous):